Consider the following. / 18-x (8 - x)2 dx (a) Perform the integration in two ways: once using the Simple Power Rule and oncePlease help with general power rule
Answer #1

| 18-aje da = f (64-16ata?) da =648-16 at ti + x2+1 + 2+1 -64a-to al t an t 3 =1648-88² +3 + c General Power Rule (18-afd da-b) Be cause ause С gepresents any constant, c-812 can be woitten Co The two answeos represent the auos family of functions Ge